OpenStudy (anonymous):

Help please!! ((: Solve each equation by factoring:

OpenStudy (anonymous):

\[x^2-11x+19=-5\]

OpenStudy (anonymous):

Haha(: wow. thanx:P

OpenStudy (anonymous):

Add 5 to both sides. x^2 - 11x + 24 = 0 (x - 8)(x - 3) = 0 So, either x -8 =0 or x-3 =0 X is both 8 and 3, written this way: 3, 8

OpenStudy (anonymous):

Haha(: It made me laugh @uri :P @jslp3z Thanks!

OpenStudy (anonymous):

No problem(: